OpenStudy (anonymous):

What is the equation of the line that passes through the points (–12, –8) and (–17, –16)?

OpenStudy (anonymous):

How? I need help!

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

First thing to do is to find the slope of the line created by those 2 points. Can you do that? Do you know the formula?

OpenStudy (anonymous):

is it, y-y1=m(x-x1)

OpenStudy (imstuck):

That's the point-slope formula. We will get there, but not til after we find the slope.

OpenStudy (anonymous):

okay so how do we find the slope

OpenStudy (imstuck):

The slope formula is this one:\[m=\frac{ y _{2}-y _{1} }{ x _{2} -x _{1}}\]

OpenStudy (anonymous):

oh i think i might be able to get it one sec

OpenStudy (imstuck):

use the points they gave you like this:

OpenStudy (imstuck):

\[m=\frac{ -16-(-8) }{ -17-(-12) }\]

OpenStudy (imstuck):

Let me know what you get, ok?

OpenStudy (anonymous):

8/5 or -8/-5

OpenStudy (imstuck):

Right! It's 8/5. Now use one of the points and that slope you just found in the point-slope equation.

OpenStudy (anonymous):

okay like the first set of points

OpenStudy (imstuck):

Want me to show you?

OpenStudy (anonymous):

is it like this, y-(-8)=8/5(x-(-12)

OpenStudy (imstuck):

yes it is! Now can you solve it? By that I mean to find the equation and put it into slope-intercept form.

OpenStudy (anonymous):

but the -8 and -12 will be positive

OpenStudy (imstuck):

Yes they will be. You're right...don't doubt yourself...you got this! Go on...

OpenStudy (imstuck):

What form does the equation have to be in for the final answer? Point-slope or slope-intercept?

OpenStudy (anonymous):

it will equal into the slope intercept form to, y=8/5x + 11.2

OpenStudy (imstuck):

That is so very right! Good! Very good!

OpenStudy (anonymous):

thanks know i know how to completely do them thanks

OpenStudy (imstuck):

You are so welcome! TY for the medal!

OpenStudy (anonymous):

Np